# Healthy LLMs? Benchmarking LLM Knowledge of UK Government Public Health Information

### v1

@@ Abstract
As Large Language Models (LLMs) become widely accessible, a detailed understanding of their knowledge within specific domains becomes necessary for successful real world use. This is particularly critical in public health, where failure to retrieve relevant, accurate, and current information could significantly impact UK residents. However, currently little is known about LLM knowledge of UK Government public health information. To address this issue, this paper introduces a new benchmark, PubHealthBench, with over 8000 questions for evaluating LLMs’ Multiple Choice Question Answering (MCQA) and free form responses to public health queries, created via an automated pipeline. We also release a new dataset of the extracted UK Government public health guidance documents used as source text for PubHealthBench. Assessing 24 LLMs on PubHealthBench we find the latest private LLMs (GPT-4.5, GPT-4.1 and o1) have a high degree of knowledge, achieving >90% in the MCQA setup, and outperform humans with cursory search engine use. However, in the free form setup we see lower performance with no model scoring >75%. Therefore, whilst there are promising signs that state of the art (SOTA) LLMs are an increasingly accurate source of public health information, additional safeguards or tools may still be needed when providing free form responses on public health topics.
@@ 1 Introduction
In this paper we introduce a new dataset, Multiple-Choice Question Answering (MCQA) benchmark, and free form response benchmark for assessing LLMs’ knowledge across a broad range of UK public health guidance [^1] The guidance included can cover the entire UK or individual constituent countries, most documents primarily relate to English public health guidance. We also only included English language documents.. Specifically our contributions include:
UK public health guidance dataset [^2] Documents: https://huggingface.co/datasets/Joshua-Harris/PubHealthBench-documents - We collect, extract, markdown format, and chunk information from over 1000 publicly available UK Health Security Agency (UKHSA) PDF and HTML documents from the UK Government website (gov.uk). We make the extracted text from a filtered dataset (687 documents) publicly available in markdown format to enable further research into the use of LLMs on UK public health advice.
PubHealthBench [^3] Benchmark: https://huggingface.co/datasets/Joshua-Harris/PubHealthBench a fully grounded MCQA benchmark - We implement an automated MCQA generation and validation pipeline grounded in the extracted guidance source text. This enables us to generate a new benchmark with over 8000 MCQA questions to test LLM knowledge across a broad range of current guidance. We provide results for both the full benchmark (PubHealthBench-Full) and a manually reviewed subset (PubHealthBench-Reviewed).
@@
Manual expert review and human baseline - To quality assure the benchmark and establish an upper bound for performance, human experts manually reviewed a random sample of 800 MCQA examples (approx. 10% of the benchmark). Furthermore, in order to compare to human performance, 5 humans also took sample tests (total 600 MCQA examples) to establish a human baseline with the use of search engines. This provides an initial indication of the boundary at which LLMs become similarly accurate to a cursory search by non-expert humans.
@@ 5.3 PubHealthBench-FreeForm results
The free form response setup is substantially more challenging for a few reasons: (1) it requires recalling the correct guidance information without any hints from MCQA options, (2) it introduces the possibility of LLMs hallucinating additional information that may be inconsistent with the source text, and (3) the correct answer cannot be inferred via elimination of the other answer options. As a result, all models achieve substantially lower scores in the free form setting by up to 60 percentage points (ppts). The best performing model (o1) scores 74% (Table 5.3) and also sees the smallest decline from its MCQA performance at -17ppts (Table 5).
@@
Unlike in the MCQA setup, we see initial indications that "reasoning" models (o1 and o3-Mini) start to outperform "non-reasoning" models . Also, importantly, as observed in the MCQA setting, LLMs consistently show more accurate knowledge of guidance intended for the general public compared to clinical or professional guidance.
@@ 8 Conclusion
We hope this new benchmark and UK Government guidance dataset will enable further research and evaluation of LLMs for public health, and reduce the risks within real-world deployments.

### v2

@@ Abstract
As Large Language Models (LLMs) become widely accessible, a detailed understanding of their knowledge within specific domains becomes necessary for successful real world use. This is particularly critical in public health, where failure to retrieve relevant, accurate, and current information could significantly impact UK residents. However, currently little is known about LLM knowledge of UK Government public health information. To address this issue, this paper introduces a new benchmark, PubHealthBench, with over 8000 questions for evaluating LLMs’ Multiple Choice Question Answering (MCQA) and free form responses to public health queries. To create PubHealthBench we extract free text from 687 current UK government guidance documents and implement an automated pipeline for generating MCQA samples. Assessing 24 LLMs on PubHealthBench we find the latest private LLMs (GPT-4.5, GPT-4.1 and o1) have a high degree of knowledge, achieving >90% accuracy in the MCQA setup, and outperform humans with cursory search engine use. However, in the free form setup we see lower performance with no model scoring >75%. Importantly we find in both setups LLMs have higher accuracy on guidance intended for the general public. Therefore, there are promising signs that state of the art (SOTA) LLMs are an increasingly accurate source of public health information, but additional safeguards or tools may still be needed when providing free form responses on public health topics.
@@ 1 Introduction
In this paper we introduce a new dataset, Multiple-Choice Question Answering (MCQA) benchmark, and free form response benchmark for assessing LLMs’ knowledge across a broad range of UK public health guidance. [^1] The guidance included can cover the entire UK or individual constituent countries, most documents primarily relate to English public health guidance. We also only included English language documents. Specifically our contributions include:
PubHealthBench [^2] Benchmark: https://huggingface.co/datasets/Joshua-Harris/PubHealthBench a fully grounded MCQA benchmark - We collect, extract, markdown format, and chunk information from over 500 publicly available UK Health Security Agency (UKHSA) PDF and HTML documents from the UK Government website (gov.uk). [^3] GOV.UK reuse policy: https://www.gov.uk/help/reuse-govuk-content We implement an automated MCQA generation and validation pipeline grounded in the extracted guidance source text. This enables us to generate a new benchmark with over 8000 MCQA questions to test LLM knowledge across a broad range of current guidance. We provide results for both the full benchmark (PubHealthBench-Full) and a manually reviewed subset (PubHealthBench-Reviewed).
@@
Manual human expert review and human baseline - To quality assure the benchmark and establish an upper bound for performance, human experts manually reviewed a random sample of 800 MCQA examples (approx. 10% of the benchmark). Furthermore, in order to compare to human performance, 5 humans also took sample tests (total 600 MCQA examples) to establish a human baseline with the use of search engines. This provides an initial indication of the boundary at which LLMs become similarly accurate to a cursory search by non-expert humans.
@@ 5.3 PubHealthBench-FreeForm results
The free form response setup is substantially more challenging for a few reasons: (1) it requires recalling the correct guidance information without any hints from MCQA options, (2) it introduces the possibility of LLMs hallucinating additional information that may be inconsistent with the source text, and (3) the correct answer cannot be inferred via elimination of the other answer options. As a result, all models achieve substantially lower scores in the free form setting by up to 60 percentage points (ppts). The best performing model (o1) scores 74% (Table 4) and also sees the smallest decline from its MCQA performance at -17ppts (Table 5).
@@
Unlike in the MCQA setup, we see initial indications that "reasoning" models (o1 and o3-Mini) start to outperform "non-reasoning" models, showing the smallest declines from the MCQA setup. Also, importantly, as observed in the MCQA setting, LLMs consistently show more accurate knowledge of guidance intended for the general public compared to clinical or professional guidance.
@@ 8 Conclusion
We hope this new benchmark will enable further research and evaluation of LLMs for public health, and reduce the risks within real-world deployments.